Fermi National Laboratory


Accelerator Update

Friday April 23
The day shift began with Operations monitoring store 3434 and a Recycler stash, with the Antiproton Source (Pbar) stacking, and with MiniBooNE and the Meson MCenter experiments taking beam.

At 8:26 AM, Operations reported that beam had stopped making it to the Main Injector (MI) due to a kicker problem. Experts investigated and found a bad amplifier.

The MCenter experimenters went on access at 9:42 AM.

A technician replaced a time delay relay for a MI quadrupole power supply at 10:11 AM.

At 10:14 AM, Linac experts made a fifteen minute access into the H- Source dome to check the source bottle.

Pbar resumed stacking at 11:03 PM.

A MI expert began a slipstacking study at 12:54 PM.

At 1:18 PM, the Collider Detector at Fermilab (CDF) called about beam problems. Operators tried changing the tune and moving in some collimators, but neither worked. CDF turned off their high voltages and stopped collecting data at 1:36 PM.

Operations resumed sending beam to MCenter at 1:21 PM.

Operators did further collimator moves that made the beam acceptable to CDF around 2:16 PM.

The MCenter experimenters went on access at 4:20 PM.

The MI expert completed her studies at 5:16 PM.

Operations halted MI beam for a Recycler study at 7:02 PM. MI resumed taking beam at 7:26 PM.

Operations interlocked the Meson MTest experiment at 8:52 PM and began sending beam at 9:07 PM.

MCenter went in and out of access many times this evening, but at 11:16 PM, they were taking beam again.

Saturday April 24
The midnight shift began with Operations monitoring store 3434 and the Recycler stash, with Pbar stacking, and with MiniBooNE, MCenter, and MTest taking beam.

The MCenter experimenters went on access at 2:46 AM.

At 2:30 AM, Operations terminated store 3434.

Operations turned off beam to Switchyard (SY) and put the Tevatron (TeV) into shot setup at 3:08 AM.

At 4:22 AM, Operations tried their first extraction of antiprotons from Pbar, but got nothing. Tried again at 4:30 AM and still nothing. Operators talked to an expert, made some changes, but had no better luck on the third shot.

The Booster suffered a vacuum burst that tripped off two RF stations (BRF8 & 9). They both reset okay.

The antiproton transfers finished at 5:51 AM. The problems with the transfers had to do with an Accumulator RF station (ARF4), with the TeV not closing on the antiprotons due to a troublesome scope, and due to MI not closing on two of the transfers.

Operations established store 3436 at 6:11 AM with an initial luminosity of 47.62E30. Pbar resumed stacking.

Operations restored beam to MTest and MCenter at 6:22 AM.

Operations setup Pbar and Recycler for stashing antiprotons at 7:54 AM. This halted all SY beam.

By 9:44 AM, Operations had stopped stashing and resumed stacking.

Operations resumed sending beam to SY at 10:07 AM.

At 3:54 PM, a TeV RF station (TRF3) tripped off due to a power supply control failure. Operators investigated and discovered an LCW leak. They valved off the station and informed an expert.

Per instructions from a TeV expert, Operations turned off the TeV longitudinal dampers at 4:43 PM.

MCenter experimenters accessed MC7 at 5:03 PM. There were numerous problems that operators and the Duty Mechanic had to resolve before Operations restored beam about three and half hours later.

An RF expert came in at 7:11 PM and worked on TRF3 and MIRF18. He had them both back on line within thirty minutes.

A Pbar expert halted stacking to conduct a reverse proton study at 9:27 PM.

A MI expert began a study at 10:12 PM.

MCenter experimenters accessed MC7 at 11:31 PM.

Sunday April 25
The midnight shift began with Operations monitoring store 3436 and the stash, with Pbar and MI conducting studies, and with MiniBooNE and MTest taking beam.

At 12:14 AM, the MI study ended.

MCenter came out of access at 1:36 AM.

The Pbar study ended and Pbar resumed stacking at 4:42 AM.

Operations received a sump failure alarm for Meson enclosure M01 at 5:31 AM. Operators turned off the beam to Meson and prepared the area for access. Operations reported at 7:00 AM that they found nothing wrong with the M01 sump and restored beam to Meson at 7:01 AM. On a later access, the Duty Mechanic fixed a float for another sump pump.

A MI expert began a two-hour barrier bucket study at 8:40 AM.

MCenter experimenters accessed MC7 at 10:52 AM.

Operations terminated store 3436 at 1:06 PM.

At 1:51 PM, while operators were putting the TeV through a pre-shot setup a device at F1 caused the TeV to abort the beam. Operators talked to an expert and by 3:34 PM had replaced a crate's power supply and a device chassis.

MCenter came out of access at 2:20 PM.

Operations put the TeV through a dry squeeze at 2:26 PM.

At 3:44 PM, Operations reported finding a bad CAMAC crate power supply. They had a new supply installed by 4:21 PM.

At 4:43 PM, Operations reported having problems with Booster extraction. They discovered a kicker (MKS07) not firing. A proton source expert investigated while operators tuned beam around the kicker.

Operations put the TeV through a dry squeeze at 4:46 PM.

MTest came out of access at 5:48 PM.

MCenter experimenters accessed MC7 at 6:41 PM.

Two Booster experts completed their repairs on MKS07 at 7:22 PM.

MCenter came out of access at 7:27 PM.

At 8 PM, operators discovered a TeV vertical separator that hadn't come up to full voltage.

Operations aborted the TeV protons and started Pbar stacking again at 8:06 PM while investigating the separator problem. Operators reported that the separator had failed during the injection tune-up. At 8:19 PM, Operations put the TeV successfully through a wet squeeze. An expert was called in to help investigate.

Operations reestablished beam to SY at 9:14 PM.

The TeV expert worried that the there might have been some arcing in the separator tank. At 10:19 PM, he asked Operations to put the TeV through another dry squeeze. At 11:04 PM, Operations put the TeV through another dry squeeze.

Operations turned of SY beam and put the TeV into shot setup 11:27 PM.

Monday April 26
The midnight shift began with the TeV in shot setup and SY off, with Operations monitoring the stash, and with MiniBooNE taking beam.

At 2:25 AM, Operations established store 3444 with an initial luminosity of 54.33E30. Pbar resumed stacking.

At 2:46 AM, Operations shutdown Linac beam due to a Klystron water temperature alarm. Operators found a pump on the Klystron RF water skid not working properly. An expert came in at 3:55 AM and investigated. Operations talked to another expert and power cycled the temperature controller and put a mixing valve in manual mode. By 4:41 AM, the temperature was back down into the normal range.

Pbar resumed stacking at 3:22 AM.

A Recycler expert began stashing beam at 4:15 AM. At 5:11 AM, he successfully completed a record stash of 134.08E10. Pbar resumed stacking.

Plans for Today
The plans for today are to stack and store. When the stack reaches a reasonable level sometime this afternoon, Pbar will conduct some studies.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/26/2004   email Fermilab